OpenStudy (anonymous):

solve the following system. y=x+4 5x+y=16

OpenStudy (anonymous):

(2,-6) (-2,6) (6,2) (2,6)

OpenStudy (ajprincess):

substitute x+4 in place of y in the second equation and solve for x

OpenStudy (ajprincess):

5x+y=16 5x+x+4=16

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

or was that wat im supposed to do?

OpenStudy (ajprincess):

hav u found x?

OpenStudy (anonymous):

i got 2?

OpenStudy (ajprincess):

ya u r right:) good work:) nw to find y plug in 2 in place of x in y=x+4

OpenStudy (anonymous):

6?

OpenStudy (ajprincess):

yup:) very good:)

OpenStudy (anonymous):

so its (2,6) ??

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

thanks :)

OpenStudy (ajprincess):

welcome:)